lionel delivers milk, bread, and eggs to marcies camp store. on monday the bill for 8 half-gallons of milk, 4 loaves of bread, and 6 dozen eggs was $14.20. on tuesday the bill for 5 half-gallons of milk, 10 loaves of bread, and 3 dozen eggs was $14.50. on wednesday the bill for 2 half-gallons of milk, 5 loaves of bread, and 7 dozen eggs was $9.50. how much is thursdays bill for 1 half-gallon of milk, 2 loaves of bread, and 1 dozen eggs?
Answer by ikleyn(52806)

.
lionel delivers milk, bread, and eggs to marcies camp store.
on monday the bill for 8 half-gallons of milk, 4 loaves of bread, and 6 dozen eggs was $14.20.
on tuesday the bill for 5 half-gallons of milk, 10 loaves of bread, and 3 dozen eggs was $14.50.
on wednesday the bill for 2 half-gallons of milk, 5 loaves of bread, and 7 dozen eggs was $9.50.
how much is thursdays bill for 1 half-gallon of milk, 2 loaves of bread, and 1 dozen eggs?
~~~~~~~~~~~~~~~~~

Let x = price of half-gallon of milk;

    y = price of one loave of bread, 

and z = price of one dozen of eggs.



Based on the problem's info we have this system of 3 equations in 3 unknowns


    8x +  4y + 6z = 14.20  dollars    (1)   (the bill on Monday)

    5x + 10y + 3z = 14.50  dollars    (2)   (the bill on Tuesday)

    2x + 5y + 7z =  9.50  dollars     (3)   (the bill on Wednesday)



Solve the system.


You may use any method you want, or you may use your calculator, or one of numerous online solvers in the Internet.



The solution is  x = 1 dollar (milk);  y = 0.8 dollar (bread);  z = 0.5 dollars (dozen of eggs).



Having it, calculate the  ANSWER  to the problem's question  1 + 2*0.8 + 0.5 = 3.10 dollars.

Solved, answered and explained.